OpenStudy (anonymous):

Factor, check by multiplying factors. 3ab-a-3b^2+b

OpenStudy (anonymous):

@JungHyunRan

OpenStudy (anonymous):

who knew that social media needed help with math homework :P

OpenStudy (anonymous):

\(\Large 3ab-a-3b^2+b=(3ab-3b^2)-(a-b)\\\Large =3b(a-b)-(a-b)=(3b-1)(a-b)\)